# Inflation with vector fields revisited: heavy entropy perturbations and primordial black holes

### v2

@@ 4.3 PBH formation
After treating the power spectrum on large and small scales for board and sharp case, we can also compare them with the amplitude of the curvature power spectrum required for PBH formation. The fraction of PBHs at formation time can be estimated in a model-independent way [59]
@@
where $\gamma$ can be calculated analytically as $\gamma=0.2$ [51], $g_{*,f}$ is the effective number of relativistic degrees of freedom at formation time, $M_{\text{PBH}}$ is the mass of PBHs at formation time and $f_{\text{PBH}}\equiv\Omega_{\text{PBH}}/\Omega_{\text{DM}}$ is the present-day fraction of PBH density.
On the other hand, the $\beta$ can be related to initial conditions of matter power spectrum through collapse theory of Newtonian gravity. The well-used model is the Press-Schechter one [60]. After assuming the Gaussian distribution of overdensity and simply taking the critical density contrast as the threshold $\delta_{c}=0.4$ [61], the $\beta$ can be obtained to
@@
where $\sigma^{2}\ll\delta_{c}^{2}$ is the variance of the distribution. Smaller $\sigma$ implies less contrast exceeding the critical value, resulting in less $\beta_{\text{PBH}}$. The variance can be calculated from primordial power spectrum as [62]
@@
where the window function $W(q,R)$ is selected to Gaussian or top-hat usually and the comoving horizon size of PBH formed is related to the mass as $R^{-1}=k_{p}$, where [63]
@@
For inflation with a triad of vector fields, where both large-scale and PBH-scale are not trivial, the enhancement at PBH scale comparing with CMB scale is
@@
Tne power spectrum at CMB scale is $\mathcal{P}_{\zeta}(k_{\text{CMB}})=2\times 10^{-9}$. For board case this is given by (4.4) and (4.7), while for sharp case, the rapid oscillation will be smoothed out by window function and we can use (4.14) and (4.19).
Figure 6: The fraction of PBHs at formation $\beta_{\text{PBH}}$ v.s. height of peaks $h$. Solid curves denote the cases of $\delta N=2$ for $h_{\text{G}}$ and $\delta N=0.4$ for $h_{\text{TH}}$, while dashed curves denote the cases of smaller $\delta N$. Orange region denotes regime that $f_{\text{PBH}}=1$ between $10^{-17}M_{\odot}<M_{\text{PBH}}<10^{-12}M_{\odot}$.
If we consider the window where the PBHs can be the candidate of dark matter: $10^{-17}M_{\odot}<M_{\text{PBH}}<10^{-12}M_{\odot}$ and $f_{\text{PBH}}=1$, from (4.20) we obtain $1.2\times 10^{-17}<\beta_{\text{PBH}}<3.7\times 10^{-15}$. Then from (4.21) the variance for PBH formation in this mass regimes is $0.0510.047<\sigma<$ 0.051. When the power spectrum is peaked at $k_{p}$, the variance is reduced to $\sigma^{2}\simeq\mathcal{P}_{\zeta}(k_{p})/5$. It’s seem that we need $\mathcal{P}_{\zeta}(k_{p})\sim 10^{-2}$ between $10^{12}\text{Mpc}^{-1}<k_{p}<10^{15}\text{Mpc}^{-1}$ to generate sufficient PBHs. However, because the profile of the peak is very sensitive to $h$ and $\delta N$ and the large-scale power spectrum also depends on these parameters. And the fraction $\beta_{\text{PBH}}$ is exponentially dependent on the power spectrum. Slightly changing profile of the peak will lead to significant change in $\beta_{\text{PBH}}$. We show how $\beta_{\text{PBH}}$ depend on parameters in Figure 6 for board Gaussian peaks and sharp top-hat peaks. We have used $R=k_{p}^{-1}$, which are the scales of peaks and can be separately obtained through (4.6) and (4.15) for Gaussian and top-hat examples. The $\beta_{\text{PBH}}$ decays rapidly for smaller parameter, i.e., the smaller height and fatness of the peaks. Also, the difference of using different window functions is also distinguishable.
The mass spectrum of PBHs are shown in Figure 7 for board Gaussian peaks and in Figure 8 for sharp top-hat peaks. We choose the parameters of these examples as following table.
Figure 7: Abundance of PBHs as dark matter for board($\delta N\gtrsim 1$) Gaussian profile of peak $h_{\text{G}}$.
Figure 8: Abundance of PBHs as dark matter for shape($\delta N\lesssim 1$) top-hat profile of peak $h_{\text{TH}}$.
@@
For solid curves, we choose the parameters so that the PBHs are responsible for total dark matter $f_{\text{PBH}}=1$ in our universe. The time of non-vanishing vector fields are chosen as $N_{f}=31.5$ for all solid examples to avoid constraints from observations of evaporation of PBHs due to Hawking radiation (EVAP) [59] and microlensing of stars in M31 by Subaru Hyper SuprimeCam (HSC) [64]. For dashed curves, we choose different parameters so that $f_{\text{PBH}}<1$ and also different $N_{f}=25.3$, which corresponds to larger PBH mass formation. We have seen in Figure 6 the abundance is sensitive to these parameter. As the first and the third column show, a mere 1% decrease in $h$ leads to a drastic drop in abundance from 1 to $10^{-5}$. For top-hat profile of peak $h_{\text{TH}}$, the abundance is also sensitive to $\delta N$ (4.14). So we can see in the last column of table, 1% decrease in $\delta N$ also reduces $f_{\text{PBH}}$ to $10^{-5}$.
We also compare the results of using Gaussian and top-hat window function. The choice of way of smoothing out the fluctuations increases the uncertainty in the results of PBH formation [65, 66]. We demonstrate a significant difference in the parameters required when using different window functions. For example, the first row in table, To generate the same abundance of PBHs, the required $h$ differs by up to 20% when using Gaussian and top-hat window functions. We usually need a larger primordial density in Gaussian smoothing. Furthermore, we also find that using different window functions results in variations in the peak position of $f_{\text{PBH}}$. The mass of formation is usually greater when using top-hat window function. In the examples, this difference reaches up to $M_{\text{peak,TH}}/M_{\text{peak,G}}\sim 10$.
@@ 5 Discussion
In addition to driving inflation, the enhancement of the power spectrum can also serve as a mechanism for the generation of primordial black holes (PBHs). This requires a strong excitation of the vector field during the late stages of inflation. We discussed two toy examples of $h$: Gaussian and top-hat. The Gaussian is employed to discuss the broad case, while the top-hat is used to address the sharp case, as both yield more precise analytical results. Due to the non-decaying entropy perturbations, the large-scale power spectrum is shifted compared to the single-field one. The amplitude of the shift depends on the details of vector fields dynamics, i.e., $\delta N$ and $h$, as shown in (4.7) and (4.19). For the shape top-hat case, this amplitude is only a function of $\delta N$, allowing for an analytical calculation of the shape of the power spectrum. Given the exponential dependence of the power spectrum shape on these parameters and the subsequent exponential dependence of PBH abundance on the power spectrum shape, the PBH abundance is highly sensitive to these parameters. Whether it’s $h$ or $\delta N$, even slight variations can significantly reduce the abundance of PBHs. Certainly, this sensitivity is also related to the specific details of the PBH collapse. For instance, we demonstrated that under the same $\delta_{c}$, smoothing the same fluctuations in different ways can lead to vastly different results, not only in PBH abundance but also in the masses of PBH formation. For example, compared to a Gaussian window function, using a top-hat window function typically results in PBHs with smaller masses. Further investigation is required to address this question comprehensively.
@@
Now, we delve into another regime, characterized by a relatively large energy of the vector field. In this scenario, the contributions to inflationary predictions differ significantly from the case where $h$ is very small. Many questions need to be reconsidered in this regime, such as the anisotropy for one vector field and non-Gaussianity. We posit that the discussions in this paper are also applicable to a single vector field. In this case, the $IN^{2}$ term in power spectrum, which served as a contribution to anisotropy for small $h$ case, is replaced by other constant modes. It is imperative for us to calculate the anisotropy in this regime. Additionally, when considering a charged field as the inflaton, whether the perturbative Schwinger effect from the electric field is crucial to inflationary predictions is also a question that warrants consideration. Regarding PBHs, we are discussing the isotropic configuration of vector fields as an example. Although it bears similarities to hyperbolic inflation, in general, vector fields themselves exhibit phenomenology distinct from scalar fields, such as anisotropy for a single vector field and chirality for symmetry breaking. These differences manifest in the generated PBHs and induced gravitational waves. Recently observations from 15-year NANOGrav data set reveals information about low-frequency stochastic gravitational wave background [67]. PBHs can be one of the sources of this background [68]. The formation of PBHs in single-field up to one-loop correction is still controversial [69, 70, 71, 72, 73, 74, 75, 76]. The single-field EFT with a imaginary $c_{s}$ should also be discussed. We intend to present these studies in the near future.
@@ Appendix A Derivation of the quadratic action
Appendix B Effective field theory of inflation
In the EFT formalism, the physical scalar d.o.f at low-energy scale is the Nambu-Goldstone (NG) boson particle coming from breaking the time translation symmetry $t\to t+\xi^{0}$ of the de Sitter space by a clock field $\phi(t)$, i.e., the homogeneous inflaton. One can choose the slice of $t$ that the inflaton has no excitation all the time, i.e. the unitary gauge. In this gauge, the only dynamical scalar d.o.f at low-energy is the scalar part of graviton $g_{ij}=a^{2}(t)e^{2\zeta}\delta_{ij}$, where the scalar $\zeta(t,\boldsymbol{x})$ is the spatial curvature of the time slice. The EFT of a single scalar field in the unitary gauge can be written as [39]
$\mathcal{L}_{\text{EFT}}=\sqrt{-g}\left[\frac{M_{\text{pl}}^{2}}{2}R+M_{\text{pl}}^{2}\dot{H}g^{00}-M_{\text{pl}}^{2}\left(3H^{2}+\dot{H}\right)+\cdots\right]+\mathcal{L}_{\text{n-u}},$ (B.1)
where the dots stand for the higher-order terms of $(1+g^{00})$ and we have ignored the extrinsic curvature fluctuation $\delta K_{\mu\nu}$. $\mathcal{L}_{\text{n-u}}$ is the non-universal parts and for $\omega^{2}\ll M^{2}+p^{2}$ [40]
$\mathcal{L}_{\text{n-u}}=\sqrt{-g}\sum_{n=2}\frac{M_{n}^{4}}{n!}\left[(1+g^{00})\frac{M^{2}}{M^{2}-\tilde{\partial}^{2}}\right]^{n-1}(1+g^{00})+\cdots,$ (B.2)
where $M$ is the mass of heavy fields and ~$\tilde{\partial}\equiv\partial/a$. These operators are non-local generally due to the hierarchy between the frequency and momentum when the speed of sound is small $c_{s}\ll 1$. There exist regime in energy scales that the dispersion is non-linear [40, 41]. If $\omega\sim p$ the momentum is also $p\ll M$ and we can ignore the non-local operators.
In order to transform to the spatial flat gauge $g_{ij}=a^{2}(t)\delta_{ij}$, one can perform a time diffeomorphisms $t\to\tilde{t}=t-\pi(t,\boldsymbol{x})$. Then at linear-order the NG boson is related to the curvature fluctuation as $\zeta=-H\pi$. The time component of the metric transforms as
$g^{00}=\frac{\partial x^{0}}{\partial\tilde{x}^{\mu}}\frac{\partial x^{0}}{\partial\tilde{x}^{\nu}}\tilde{g}^{\mu\nu}=\left(1+\dot{\pi}\right)^{2}\tilde{g}^{00}+2\left(1+\dot{\pi}\right)\partial_{i}\pi\tilde{g}^{0i}+\tilde{g}^{ij}\partial_{i}\pi\partial_{j}\pi$ (B.3)
In the low-energy scale $\omega^{2}\ll p^{2}+M^{2}$, we can write down the action in the re-parameterization of time. Dropping the tildes for simplicity, the Lagrangian of NG boson is given by [40]
$\mathcal{L}_{\pi}=$ $\sqrt{-g}\Bigg{\{}-M_{\text{pl}}^{2}\dot{H}\left[\dot{\pi}\left(1+\frac{2M_{2}^{4}}{M_{\text{pl}}^{2}|\dot{H}|}\frac{M^{2}}{M^{2}-\tilde{\partial}^{2}}\right)\dot{\pi}-(\tilde{\partial}\pi)^{2}\right]$
$+2M_{2}^{4}\left[\dot{\pi}^{2}-(\tilde{\partial}\pi)^{2}\right]\frac{M^{2}}{M^{2}-\tilde{\partial}^{2}}\dot{\pi}-\frac{4}{3}M_{3}^{4}\left(\dot{\pi}\frac{M^{2}}{M^{2}-\tilde{\partial}^{2}}\right)^{2}\dot{\pi}+\cdots\Bigg{\}},$ (B.4)
where we have expanding the action up to third order of fluctuations.
In the decoupling limit, the NG boson is decoupled with gravity and becomes a dynamical scalar mode. For the inflation, this decoupling limit is the slow-roll limit $\epsilon\to 0$. Then the quadratic Lagrangian of NG boson reduces to
$\mathcal{L}_{\pi}=a^{3}M_{\text{pl}}^{2}\epsilon\left[\dot{\pi}\left(\frac{M^{2}c_{s}^{-2}-\tilde{\partial}^{2}}{M^{2}-\tilde{\partial}^{2}}\right)\dot{\pi}-(\tilde{\partial}\mathcal{\pi})^{2}\right],$ (B.5)
where the speed of sound is defined by
$\frac{1}{c_{s}^{2}}\equiv 1-\frac{2M_{2}^{4}}{M_{\text{pl}}^{2}\dot{H}}.$ (B.6)
The modification of the speed of sound is provided by the higher operator $M_{2}$ and we will show soon that this can be obtained by integrating out the heavy fluctuation at low energy in the dilaton-gauge model. The dispersion relation of the propagation of the light field $\pi$ should be written as
$\omega^{2}(p)=\frac{M^{2}+p^{2}}{M^{2}c_{s}^{-2}+p^{2}}p^{2},$ (B.7)
which is generally non-linear. However, if the speed of sound $c_{s}$ is not very small, The region with nonlinear dispersion relations is very narrow so the Lagrangian of NG boson shortly reduces to
$\mathcal{L}_{\pi}=a^{3}M_{\text{pl}}^{2}H^{2}\frac{\epsilon}{c_{s}^{2}}\left[\dot{\pi}^{2}-\frac{c_{s}^{2}}{a^{2}}(\partial\mathcal{\pi})^{2}\right],$ (B.8)
where the dispersion relation is almost linear $\omega(p)\simeq c_{s}p$ throughout the entire process.

### v3

@@ 4.3 PBH formation
After treating the power spectrum on large and small scales for the broad and sharp case, we can also compare them with the amplitude of the curvature power spectrum required for PBH formation. The fraction of PBHs at the formation time can be estimated in a model-independent way [62]
@@
where $\gamma$ can be calculated analytically as $\gamma=0.2$ [54], $g_{*,f}$ is the effective number of relativistic degrees of freedom at formation time, $M_{\text{PBH}}$ is the mass of PBHs at formation time, and $f_{\text{PBH}}\equiv\Omega_{\text{PBH}}/\Omega_{\text{DM}}$ is the present-day fraction of PBH density.
On the other hand, the $\beta$ can be related to the initial conditions of the matter power spectrum through the collapse theory of Newtonian gravity. Acommonly used model is the Press-Schechter one [63]. After assuming a Gaussian distribution of overdensity and taking the critical density contrast as the threshold $\delta_{c}=0.4$ [64], $\beta$ can be obtained as
@@
where $\sigma^{2}\ll\delta_{c}^{2}$ is the variance of the distribution. A smaller $\sigma$ implies fewer contrasts exceeding the critical value, resulting in a smaller $\beta_{\text{PBH}}$. The variance can be calculated from the primordial power spectrum as [65]
@@
where the window function $W(q,R)$ is usually selected as Gaussian or top-hat, and the comoving horizon size of the PBH formed is related to the mass as $R^{-1}=k_{p}$, where [66]
@@
For inflation with a triad of vector fields, where both large-scale and PBH-scale are not trivial, the enhancement at the PBH scale comparing to the CMB scale is
@@
The power spectrum at the CMB scale is $\mathcal{P}_{\zeta}(k_{\text{CMB}})=2\times 10^{-9}$. For the broad case, this is given by (4.4) and (4.7), while for the sharp case, the rapid oscillation will be smoothed out by the window function and we can use (4.14) and (4.19).
Figure 7: The fraction of PBHs at formation $\beta_{\text{PBH}}$ v.s. the height of the peaks $h$. The solid curves denote the cases of $\delta N=2$ for $h_{\text{G}}$ and $\delta N=0.4$ for $h_{\text{TH}}$, while the dashed curves denote the cases of smaller $\delta N$. THe orange region denotes the regime where $f_{\text{PBH}}=1$ between $10^{-17}M_{\odot}<M_{\text{PBH}}<10^{-12}M_{\odot}$.
If we consider the window where the PBHs can be the a candidate of dark matter: $10^{-17}M_{\odot}<M_{\text{PBH}}<10^{-12}M_{\odot}$ and $f_{\text{PBH}}=1$, from (4.20) we obtain $1.2\times 10^{-17}<\beta_{\text{PBH}}<3.7\times 10^{-15}$. Then, from (4.21) the variance for PBH formation in this mass regimes is $0.0510.047<\sigma<$ 0.051. When the power spectrum is peaked at $k_{p}$, the variance is reduced to $\sigma^{2}\simeq\mathcal{P}_{\zeta}(k_{p})/5$. It seems that we need $\mathcal{P}_{\zeta}(k_{p})\sim 10^{-2}$ between $10^{12}\text{Mpc}^{-1}<k_{p}<10^{15}\text{Mpc}^{-1}$ to generate sufficient PBHs. However, because the profile of the peak is very sensitive to $h$ and $\delta N$, and the large-scale power spectrum also depends on these parameters. The fraction $\beta_{\text{PBH}}$ is exponentially dependent on the power spectrum. Slightly changing the profile of the peak will lead to a significant change in $\beta_{\text{PBH}}$. We show how $\beta_{\text{PBH}}$ depends on parameters in Figure 7 for broad Gaussian peaks and sharp top-hat peaks. We have used $R=k_{p}^{-1}$, which are the scales of the peaks and can be separately obtained through (4.6) and (4.15) for Gaussian and top-hat examples. The $\beta_{\text{PBH}}$ decays rapidly for smaller parameters, i.e., smaller height and fatness of the peaks. Also, the difference in using different window functions is also distinguishable.
Figure 8: Abundance of PBHs as dark matter for broad($\delta N\gtrsim 1$) Gaussian profile of peak $h_{\text{G}}$.
Figure 9: Abundance of PBHs as dark matter for sharp($\delta N\lesssim 1$) top-hat profile of peak $h_{\text{TH}}$.
The mass spectrum of PBHs are shown in Figure 8 for broad Gaussian peaks and in Figure 9 for sharp top-hat peaks. We choose the parameters of these examples as following table.
@@
For the solid curves, we choose the parameters so that the PBHs are responsible for the total dark matter $f_{\text{PBH}}=1$ in our universe. The time of non-vanishing vector fields are chosen as $N_{f}=31.5$ for all solid examples to avoid constraints from observations of the evaporation of PBHs due to Hawking radiation (EVAP) [62] and microlensing of stars in M31 by Subaru Hyper SuprimeCam (HSC) [67]. For the dashed curves, we choose different parameters so that $f_{\text{PBH}}<1$ and also different $N_{f}=25.3$, which corresponds to the formation of larger PBH mass. We have seen in Figure 7 the abundance is sensitive to these parameter. As the first and the third column show, a mere 1% decrease in $h$ leads to a drastic drop in abundance from 1 to $10^{-5}$. For the top-hat profile of the peak $h_{\text{TH}}$, the abundance is also sensitive to $\delta N$ (4.14). So we can see in the last column of table, 1% decrease in $\delta N$ also reduces $f_{\text{PBH}}$ to $10^{-5}$.
We also compare the results of using Gaussian and top-hat window functions. The choice of the method of smoothing out the fluctuations increases the uncertainty in the results of PBH formation [68, 69]. We demonstrate a significant difference in the parameters required when using different window functions. For example, in the first row of the table, to generate the same abundance of PBHs, the required $h$ differs by up to 20% when using Gaussian and top-hat window functions. We usually need a larger primordial density in Gaussian smoothing. Furthermore, we also find that using different window functions results in variations in the peak position of $f_{\text{PBH}}$. The mass of formation is usually greater when using the top-hat window function. In these examples, this difference reaches up to $M_{\text{peak,TH}}/M_{\text{peak,G}}\sim 10$.
@@ 5 Discussion
In addition to driving inflation, the enhancement of the power spectrum can also serve as a mechanism for the generation of primordial black holes (PBHs). This requires a strong excitation of the vector field during the late stages of inflation. We discussed two toy examples of $h$: Gaussian and top-hat. The Gaussian is employed to discuss the broad case, while the top-hat is used to address the sharp case, as both yield more precise analytical results. Due to the non-decaying entropy perturbations, the large-scale power spectrum is shifted compared to the single-field case. The amplitude of this shift depends on the dynamics of the vector fields, i.e., $\delta N$ and $h$, as shown in (4.7) and (4.19). For the sharp top-hat case, this amplitude is only a function of $\delta N$, allowing for an analytical calculation of the shape of the power spectrum. Given the exponential dependence of the power spectrum’s shape on these parameters and the subsequent exponential dependence of PBH abundance on the power spectrum’s shape, the PBH abundance is highly sensitive to these parameters. Whether it’s $h$ or $\delta N$, even slight variations can significantly reduce the abundance of PBHs. Certainly, this sensitivity is also related to the specific details of PBH collapse. For instance, we demonstrated that, under the same $\delta_{c}$, smoothing the same fluctuations in different ways can lead to vastly different results, not only in PBH abundance but also in the masses of the PBH formation. For example, compared to a Gaussian window function, using a top-hat window function typically results in PBHs with smaller masses. Further investigation is required to address this question comprehensively.
@@
Now, we delve into another regime, characterized by a relatively large energy of the vector field. In this scenario, the contributions to inflationary predictions differ significantly from the case where $h$ is very small. Many questions need to be reconsidered in this regime, such as the anisotropy for a single vector field, and non-Gaussianity. We posit that the discussions in this paper are also applicable to a single vector field. In this case, the $IN^{2}$ term in the power spectrum, which served as a contribution to anisotropy in the small $h$ case, is replaced by other constant modes. It is imperative for us to calculate the anisotropy in this regime. Additionally, when considering a charged field as the inflaton, whether the perturbative Schwinger effect from the electric field is crucial to inflationary predictions is also a question that warrants consideration. Regarding PBHs, we focus on the isotropic configuration of the vector fields as an example. Although it bears similarities to hyperbolic inflation, in general, vector fields themselves exhibit phenomenology distinct from scalar fields, such as anisotropy for a single vector field and chirality for symmetry breaking. These differences manifest in the generated PBHs and the induced gravitational waves. Recently observations from 15-year NANOGrav data set reveals information about low-frequency stochastic gravitational wave background [70]. PBHs can be one of the sources of this background [71]. The formation of PBHs in single-field up to one-loop correction is still controversial [72, 73, 74, 75, 76, 77, 78, 79]. The single-field EFT with a imaginary $c_{s}$ should also be discussed. We intend to present these studies in the near future.
@@ Appendix A Derivation of the quadratic action
Appendix B Modified dispersion relation
$\ddot{\zeta_{c}}+p^{2}\zeta_{c}=$ $\sqrt{2}hH\mathcal{A}\dot{\mathcal{F}},$
¨$\ddot{\mathcal{F}}+p^{2}\mathcal{F}+m_{s}^{2}\mathcal{F}=$ $-\sqrt{2}hH\mathcal{A}\dot{\zeta}_{c},$ (B.1)
$\zeta_{c}=\zeta_{+}e^{i\omega_{+}t}+\zeta_{-}e^{i\omega_{-}t},$
$\mathcal{F}=\mathcal{F}_{+}e^{i\omega_{+}t}+\mathcal{F}_{-}e^{i\omega_{-}t},$ (B.2)
where the two frequencies $\omega_{\pm}$ are given by
$\omega_{\pm}^{2}=\frac{m_{s}^{2}}{2c_{s}^{2}}+p^{2}\pm\frac{m_{s}^{2}}{2c_{s}^{2}}\sqrt{1+\frac{4\left(1-c_{s}^{2}\right)p^{2}}{m_{s}^{2}c_{s}^{-2}}},$ (B.3)
where the speed of sound is given by
$\frac{1}{c_{s}^{2}}=1+\frac{32H^{2}h^{2}}{m_{s}^{2}}.$ (B.4)
The speed of sound can be given by the dispersion relation of low-frequency modes at low-energy $4(1-c_{s}^{2})p^{2}\ll m_{s}^{2}c_{s}^{-2}$
$\omega_{-}^{2}\simeq c_{s}^{2}p^{2}.$ (B.5)
The modification of the dispersion relation of fluctuations is provided by the interaction with $\dot{\zeta}{\mathcal{F}}$. The heavy entropy fluctuation implies the hierarchy of the two frequencies $\omega_{-}\ll\omega_{+}$. Hence we have a cut-off scale on momentum
$p_{\text{UV}}^{2}=\frac{|m_{s}^{2}c_{s}^{-2}|}{4\left(1-c_{s}^{2}\right)}.$ (B.6)